# Principled Design of Interpretable Automated Scoring for Large-Scale Educational Assessments

### v1

@@ 4.1 Scoring Accuracy Experiment
The following baseline models were compared :
@@
The following LLM-based classifiers were fine-tuned on the response-score pairs: BERT Devlin et al. (2019), DeBERTa He et al. (2020), Llama-3.1-8b, and Llama-3.1-8b-Instruct Grattafiori et al. (2024). We also fine-tune Llama-3.1-8B-Instruct with a rubric of the score categories added to the input.
@@
Methods included are: AutoSAS Kumar et al. (2019), AsRRN Li et al. (2023), and NAM Condor and Pardos (2024).
@@
4.2 Featurization Alignment Experiment
To answer these questions, we sampled 50 (response, analytic component) pairs for each of the 3 assessment areas. To ensure balanced representation, the sample included a balanced number of pairs from each of the three score categories, as initially determined by the GPT-4.1-mini featurizer. We then asked 7 human annotators to conduct the labeling task on these samples. The human annotators consisted of five volunteers from an R1 university and two of the study’s authors. None of the annotators had prior exposure to any of the LLM’s featurization outputs. All annotators had advanced academic training (PhD-level) and teaching experience, five of whom have been instructors at the primary, secondary, and/or post-secondary level.
The annotators received an oral presentation of the purpose of the study along with links to 3 Qualtrics forms to be filled out, one in each assessment area. The form reiterated the study’s purpose, explained the task, and presented 50 items to annotate, each containing the context of the assessment item and the same featurizer prompt shown in Figure 2. The overall process took each annotator between 2.5 to 3.5 hours.
5 Experiment Results
5.1 Scoring Accuracy Results
Table 2 shows the results of the scoring accuracy experiments. Across items and within each assessment area, AnalyticScore outperforms111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. several automated scoring baselines on average and, given its interpretability, achieves reasonable performance compared to state-of-the-art black-box models. Except for the untuned Llama featurizer, each AnalyticScore variant outperforms111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. the few-shot prompting and automated scoring baselines. Compared to the best-performing models in each assessment area, these three AnalyticScore models are, on average, within 0.06 QWK over all items, 0.04 QWK for Science, 0.08 QWK for Reading (Informational Text), and 0.09 QWK for Reading (Literature) items.
@@ 5.1 Scoring Accuracy Results
Table 2: Test-time Quadratic Weighted Kappa (QWK) of scoring models per item, along with average per assessment area. Best, second-best, and at human-level performance scores are marked respectively. Sci.: Science (Q1,2,5,6). R(Inf): Reading(Informational Text) (Q3,4,9). R(Lit): Reading(Literature) (Q7,8). (*) are methods that are considered interpretable.
5.2 Featurization Alignment Results
[…]
Next, alignment between each featurizing model with the human ratings using majority vote is shown in Table 3. Most notably, the distilled Llama featurizer achieves substantially high agreement with the aggregate human features across all assessment areas. Other featurizers also achieve high agreement in Science and Reading (Information Text) but achieves moderate agreement in Reading (Literature).
Assessment Area Featurizer Model QWK Label Distribution6 Label-wise F1
@@ 5.2 Featurization Alignment Results
Reading
(Informational Text)
Human 11.64% 18.53% 69.83%
GPT-4.1-mini (0.72, 0.72) 9.82% 22.35% 67.83% (0.68, 0.69) (0.54, 0.55) (0.87, 0.88)
o4-mini (0.81, 0.81) 18.30% 16.60% 65.10% (0.73, 0.74) (0.68, 0.69) (0.94, 0.94)
Llama-3.1-8B (Distilled) (0.72, 0.73) 20.51% 10.16% 69.34% (0.61, 0.62) (0.24, 0.26) (0.91, 0.91)
Reading
Human 9.48% 6.57% 83.95%
GPT-4.1-mini (0.54, 0.56) 2.60% 13.86% 83.54% (0.45, 0.47) (0.67, 0.69) (0.95, 0.95)
o4-mini (0.52, 0.54) 7.22% 6.80% 85.98% (0.50, 0.52) (0.12, 0.14) (0.92, 0.92)
Llama-3.1-8B (Distilled) (0.81, 0.81) 7.22% 7.52% 85.26% (0.83, 0.84) (0.20, 0.22) (0.92, 0.92)
@@
Reading (Literature) $(0.669,0.678)$
Table 4: Inter-rater reliability among human raters for the featurization alignment experiment (95% bootstrap CI).
@@
6 Discussions and Limitations
The principles outlined in Section 2.2 address a specific aspect of “interpretability” in the broader domain of automated scoring. While these principles are foundational to supporting the design of a valid scoring process and fostering trust in the scoring system, simply adhering to the principles does not, on its own, guarantee that the needs and potential benefits of the stakeholders (Section 2.1) will be met.
Educational assessment literature is ripe with guidelines, frameworks, and best practices for ensuring that automated scoring properly serves the stakeholders’ needs and produce valid, reliable, and fair scoring results (e.g., Bennett and Bejar (1998), Bennett and Zhang (2015), Williamson et al. (2012), Bejar et al. (2016)). These studies emphasize that evidence for the validity of automated scoring should be collected throughout the assessment process from a variety of evidentiary sources, such as model features, agreement with human raters, treatment of unusual responses, generalizability of score interpretations, population invariance of scores, and impact on teaching and learning Bennett and Zhang (2015).
@@ 7 Conclusion
AI and education research community has yet to produce a practical interpretability solution for automated scoring in large-scale educational assessments despite a pressing need. To address this challenge, we analyzed the needs and potential benefits of interpretable automated scoring for various assessment stakeholders (students, assessment developers, and test users) and developed 4 foundational principles – Faithful, Grounded, Traceable, and Interchangeable (FGTI) – aimed at addressing those needs. We also demonstrated the feasibility of implementing these principles by developing AnalyticScore for short-answer scoring. This framework generates human-interpretable features for each response based on explicitly identifiable elements, and uses an intuitive ordinal logistic regression scorer. On a real-world short-answer scoring dataset, AnalyticScore outperforms many uninterpretable scoring methods, achieves a narrow performance gap relative to the uninterpretable SOTA, and demonstrates featurization behaviors that align with human judgment. Our findings show strong promise for implementing accurate and well-aligned interpretability solutions that address the real needs of assessment stakeholders. We hope our work illuminates exciting new directions in developing practical and effective interpretable automated scoring for large-scale educational assessments.

### v2

@@ 4.1 Scoring Accuracy Experiment
We compare against the following baseline methods (See Appendix C for the list of hyperparameters used):
@@
The following LLM-based classifiers were fine-tuned on the response-score pairs: BERT Devlin et al. (2019), DeBERTa He et al. (2020), Llama-3.1-8b, and Llama-3.1-8b-Instruct Grattafiori et al. (2024). We also fine-tune Llama-3.1-8B-Instruct with the scoring rubric added as a part of the input.
@@
AutoSAS Kumar et al. (2019), AsRRN Li et al. (2023), and NAM Condor and Pardos (2024).
@@
Table 1: Test-time Quadratic Weighted Kappa (QWK) of scoring models per item, along with average per assessment area. Best, second-best, and at human-level performance scores are marked respectively. Sci.: Science (Q1,2,5,6). R(Inf): Reading(Informational Text) (Q3,4,9). R(Lit): Reading(Literature) (Q7,8). (*) are methods that are considered interpretable. Human scoring accuracy was taken from Shermis (2015).
4.2 Featurization Alignment Experiment
To answer these questions, we sampled 50 (response, analytic component) pairs for each of the 3 assessment areas. To ensure balanced representation, the sample included a balanced number of pairs from each of the three score categories, as initially determined by the GPT-4.1-mini featurizer. We then asked 7 human annotators to conduct the labeling task on these samples. The human annotators consisted of five volunteers from [ANONYMOUS] and two of the study’s authors. None of the annotators had prior exposure to the LLM’s featurization outputs, preventing any potential bias. All annotators had advanced academic training (PhD-level) and teaching experience, five of whom have been instructors at the primary, secondary, and/or post-secondary level. Additional details about the annotation process can be found in Appendix D.
5 Experiment Results
5.1 Scoring Accuracy Results
Table 1 shows the results of the scoring accuracy experiments. Across items and within each assessment area, AnalyticScore outperforms111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. several automated scoring baselines on average and, given its interpretability, achieves reasonable performance compared to state-of-the-art black-box models. Except for the untuned Llama featurizer, each AnalyticScore variant outperforms111p<0.05p<0.05 for Wilcoxon signed-rank test across all items. Due to small $n$, no area-specific difference was statistically significant. the few-shot prompting and automated scoring baselines. Compared to the best-performing models in each assessment area, AnalyticScore is, on average, within 0.06 QWK for all items, 0.04 QWK for Science, 0.08 QWK for Reading (Informational Text), and 0.09 QWK for Reading (Literature) items. These results are noteworthy given that the FGTI principles strictly limit architectural flexibilities that often improve performance, such as subroutines that cannot be mapped to discrete reasoning steps or scoring components that are only partially articulable.
Assessment Area Featurizer Model QWK Label Distribution4 Label-wise F1
@@ 5.1 Scoring Accuracy Results
Reading (Informational Text) Human 11.48% 18.44% 70.08%
GPT-4.1-mini (0.71, 0.72) 9.70% 22.01% 68.29% (0.68, 0.69) (0.54, 0.55) (0.87, 0.88)
o4-mini (0.81, 0.82) 18.05% 16.54% 65.42% (0.73, 0.74) (0.69, 0.70) (0.94, 0.94)
Llama-3.1-8B (Distilled) (0.72, 0.73) 20.20% 9.98% 69.82% (0.61, 0.62) (0.24, 0.26) (0.91, 0.91)
Reading (Literature) Human 9.33% 5.12% 85.55%
GPT-4.1-mini (0.52, 0.54) 2.64% 10.67% 86.69% (0.44, 0.46) (0.67, 0.69) (0.95, 0.95)
o4-mini (0.49, 0.51) 6.16% 6.64% 87.20% (0.49, 0.51) (0.10, 0.12) (0.92, 0.92)
Llama-3.1-8B (Distilled) (0.81, 0.82) 7.45% 6.39% 86.16% (0.86, 0.87) (0.17, 0.19) (0.92, 0.92)
@@
Reading (Literature) $(0.672,0.680)$
Table 3: Inter-rater reliability among annotators for the featurization alignment experiment (95% bootstrap CI).
5.2 Featurization Alignment Results
Table 3 displays the Krippendorff’s $\alpha$ measured among the human raters in conducting the featurization task from Section 3.2. For all assessment areas, we observe $0.667\leq\alpha<0.8$, which fall into a range of acceptable inter-rater reliability Krippendorff (2018). We interpret this as a good level of rater agreement on the featurization process defined in this work and acknowledge that there is still potential to refine and improve the task further.
Next, Table 2 shows alignment between models and majority human labels. Most notably, the distilled Llama featurizer achieves substantially high agreement with the aggregate human features across all assessment areas. Other featurizers also achieve high agreement in Science and Reading (Information Text) but achieve moderate agreement in Reading (Literature).
@@ 5.2 Featurization Alignment Results
6 Discussion
6.1 Limitations and Future Work
Capturing Complex Scoring Logic
While the average performance gap of 0.06 QWK between AnalyticScore and the uninterpretable SOTA is meaningful given the architectural constraints imposed by the FGTI principles, there is still room to improve. One potential area of improvement is the choice of the scoring module used during Phase 3 (Section 3.3). Our demonstration of AnalyticScore uses ordinal logistic regression as the traceable and interchangeable scoring module. While this implementation provides a strong baseline, it may not sufficiently capture the logical nuances required for scoring complex items. Future work should explore more complex yet traceable and interchangeable alternatives to ordinal logistic regression. Examples of possible alternatives are LLM workflows or agents which consist of modules that each compute a specific evidentiary reasoning step and human-understandable intermediate outputs.
Limited Benchmark Datasets
The ASAP-SAS dataset is the largest publicly available dataset of complex open-ended short-answer scoring, but it only reflects a tiny fraction of assessment items, response types, domains, test-taker population, and assessment language. For empirical results to robustly generalize, automated scoring methods should be comprehensively evaluated on various datasets collected across different assessment settings. Given the limited amount of publicly available large-scale benchmarks for complex open-ended short-answer scoring, curation of additional datasets is necessary to support a more rigorous study of automated scoring.
Real-World Validation
Our analysis of assessment stakeholder needs and interpretability principles was scoped to a theoretical study derived from the educational assessment literature. To identify and address the gaps that exist between theory and practice, field studies should be conducted in the future to validate the needs of the real-world assessment stakeholders and iteratively refine the design of interpretable scoring systems.
Measuring Alignment
Our featurization alignment study (Sections 4.2 and 5.2) measures the model-to-human and human-to-human alignment in featurization behaviors. In addition to measuring alignment in featurization behaviors, it is also important to measure how the extracted analytic components and the featurization task align with the target constructs that the assessment item is intended to capture Bejar et al. (2016). Analyzing construct alignment requires the knowledge of the design decisions involved in the development of the assessment items, which we did not have access to in our current study. We hope to see larger-scale studies on both featurization alignment and construct alignment in authentic assessment environments in the future.
6.2 Ethical Considerations
Student Privacy
Our demonstration of AnalyticScore involves the use of proprietary LLMs for extracting analytic components, featurizing responses, and creating the training dataset to distill the featurizer. When handling real-world assessment responses, care should be taken to avoid leakage of potentially sensitive or personally identifiable information contained in student responses.
Responsible Use of Model Interpretations
To ensure that the assessment outcomes are valid, reliable, and fair, the use of automated scoring should be accompanied by the analysis of multiple sources of validity evidence Bennett and Bejar (1998), Bennett and Zhang (2015), Williamson et al. (2012), Bejar et al. (2016). These sources include the features used by the scoring model, the model’s alignment with human judgments, how the model handles unusual or unexpected responses, and the consistency of scoring behavior across student populations Bennett and Zhang (2015).
Model interpretations are a means to an end for enabling these analyses, but having an interpretable scoring system does not, on its own, ensure that the assessment outcomes will be valid, reliable, and fair. Without a thoughtfully designed operational practice around the use of model interpretations, these systems have the risk of creating a false sense of validity and trustworthiness when no proper auditing or mitigation is actually taking place. Beyond the technical and architectural aspects of interpretable automated scoring discussed in this work, it is also important to develop rigorous workflows, protocols, and systematic practices for using model interpretations to ensure that automated scoring systems are responsibly audited, monitored, and improved over time.
@@ 7 Conclusion
Despite a pressing need, the AI and education research community has yet to develop a practical interpretability framework for automated scoring in large-scale educational assessments. In this work, we presented a principled approach to address this challenge. We analyzed the needs and potential benefits of assessment stakeholders and developed four foundational principles of interpretable automated scoring. As a baseline framework for future research, we developed AnalyticScore for short-answer scoring, which achieves promising scoring accuracy and demonstrates featurization behaviors that align with human judgment. We hope this work illuminates exciting new directions in developing practical and effective interpretable automated scoring methods for large-scale educational assessments.
Appendix A Prompts
Figure 2 shows the exact prompts used to implement the feature labeling function from Section 3.2.
Figure 2: Prompts used in AnalyticScore
Appendix B ASAP-SAS Dataset Detail
Table 4 shows the details of the ASAP-SAS dataset.
Token Len. Train Valid Test Assessment Area
Q1 $47.5\pm 22.2$ 1,341 331 557 Science
Q2 $59.2\pm 22.6$ 1,024 254 426
Q3 $47.9\pm 14.6$ 1,445 363 406 Reading (Informational Text)
Q4 $40.3\pm 15.5$ 1,308 349 295
Q5 $25.1\pm 21.5$ 1,459 336 598 Science
Q6 $23.8\pm 22.6$ 1,418 379 599
Q7 $41.3\pm 25.1$ 1,432 367 599 Reading (Literature)
Q8 $53.0\pm 32.6$ 1,446 353 599
Q9 $49.7\pm 36.3$ 1,453 345 599 Reading (Informational Text)
Q10 $41.1\pm 28.5$ 1,314 326 546 Science
Table 4: ASAP-SAS dataset detail by item.
Appendix C Implementation Details
ANALYTICSCORE Implementation Details
We distilled the Llama featurizer for 2 epochs using a batch size of 4 and learning rate of 1e-4. All model calls were made through the official OpenAI API. Fine-tuning was conducted on an Ubuntu 20.04 machine with 2 RTX A6000 GPUs (49Gb memory), 16 AMD EPYC 9224 24-Core Processors, and 250Gb of CPU RAM.
Baseline Implementation Details
Below is the list of hyperparameters used in this work. Lists indicate the range of hyperparameters searched.
  • Featurizer Distillation
    Seed: 42; Learning Rate: 1e-4; Epochs: 2; Batch Size: 8; Weight Decay: 0.01; Quantization: 4-bit nf4; LoRA Rank: 16; LoRA $\alpha$: 8; LoRA Dropout: 0.05
  • Few-Shot LLM
    Number of Fewshot examples: 10 random few-shot examples per score category, sampled for each response; Temperature: 0.7
  • SFT LLM (Encoder-Only Models)
    Seed: 42; Learning rate: 3e-5; Batch Size: 12; Early Stopping: Threshold 1e-3, patience 10; Classifier Head: 2-layer feedforward network with hidden dim 32, dropout probability 0.1
  • SFT LLM (Decoder-Only Models)
    Seed: 42; Learning Rate: 1e-4; Batch Size: 8; Epochs: 20; Early Stopping: Threshold 1e-3, Patience 5; Weight Decay: 0.01; Quantization: 8-bit; LoRA Rank: 16; LoRA $\alpha$: 8; LoRA Dropout: 0.05
  • AutoSAS
    Maximum Depth: [50, 75, 100, 150, 200]; Number of Estimators: [50, 75, 100, 150, 200]; Hyperparameter selection criteria: Best QWK on validation set
    Final selected hyperparameters:
    Q1.
    Max Depth: 200; # of Estimators: 150
    Q2.
    Max Depth: 200; # of Estimators: 75
    Q3.
    Max Depth: 75; # of Estimators: 100
    Q4.
    Max Depth: 75; # of Estimators: 50
    Q5.
    Max Depth: 150; # of Estimators: 100
    Q6.
    Max Depth: 50; # of Estimators: 150
    Q7.
    Max Depth: 200; # of Estimators: 150
    Q8.
    Max Depth: 200; # of Estimators: 150
    Q9.
    Max Depth: 150; # of Estimators: 200
    Q10.
    Max Depth: 200; # of Estimators: 50
  • AsRRN
    Parameters were mostly adopted from the following public implementation: https://github.com/nkzhlee/AsRRN
    Seed: 23; Batch size: 1; Learning rate: 1e-5; Epochs: 23; LR gamma: 0.1; LR step: 15; Max sequence length: 128; Validation split: 0.2; Weight decay: 0.0001; Max gradient norm: 1.0; Warmup steps: 0.2; Gradient accumulation steps: 1; Pretraining steps: 5; Hidden dimension: 768; Message dimension: 128; Hidden dropout probability: 0.2; Number of graph steps: 2; Contrastive loss lambda: 0.01; Contrastive temperature: 1; Norm: 1
[…]
Appendix D Alignment Study Details
The annotators received an oral presentation of the purpose of the study along with links to 3 Qualtrics forms to be filled out, one in each assessment area. The form reiterated the study’s purpose, explained the task, and presented 50 items to annotate, each containing the context of the assessment item and the same featurizer prompt shown in Figure 2. The overall process took each annotator between 2.5 to 3.5 hours. Figures 3 through 5 are example screenshots from one of the three qualtrics forms used for annotation.
Figure 3: Welcome page of one of the 3 Qualtrics forms used for the featurization alignment experiment. This page contains explanations about the purpose of the study, the necessary contexts, and task description.
Figure 4: Example description of the assessment item and assessment instruction shown to the annotators.
Figure 5: Each annotation task presented annotators with a (response, analytic component) pair and asked them to select one of the three label options used in AnalyticScore’s labeling function.